OpenStudy (anonymous):

please help!! What is the volume of a sphere if it has a radius of 5 inches? (Use 3.14 for π.)

OpenStudy (anonymous):

The volume of a sphere is - \[V_s=(4/3)(\pi)(r^{3})=(4/3)(3.14)(5^3)=(4/3)(125)(3.14)=523.6 inches^3\]

OpenStudy (campbell_st):

the formula is \[V = \frac{4}{3} \pi r^3\] you need to substitute r = 5 and pi = 3.14 then evaluate

OpenStudy (anonymous):

ok, thank you so much!!